OpenStudy (anonymous):

f(t)= 1+.9e^(-.02t) with t ≥ 0 find the limit as t approaches infinity of f(t)

OpenStudy (n00bstyle):

Have you already filled in +infinity?

OpenStudy (n00bstyle):

Do you understand what happens to a graph when an exponent is negative?

OpenStudy (anonymous):

no? haha sorry what happens?

OpenStudy (n00bstyle):

okay, when an exponent is negative, and you fill in very big numbers, this 'part' of the equation will become zero

OpenStudy (n00bstyle):

You have yourself two parts: 1 and .9e^(-0.2t)

OpenStudy (anonymous):

so the .9e^(-.02t) part becomes zero? I'm not really sure what you mean

OpenStudy (n00bstyle):

Yes, that is correct, but let me rewrite it! Then it will become clear to you too!!

OpenStudy (n00bstyle):

I will give you an example: \[2^{-2}=\frac{ 1 }{ 2^{2} }=\frac{ 1 }{ 4 }\]

OpenStudy (n00bstyle):

Do you understand this?

OpenStudy (anonymous):

yes i see that

OpenStudy (n00bstyle):

okay, Now fill in an infinity, you will have instead of -2, -infinity

OpenStudy (n00bstyle):

So then again, apply the 'fraction-rule'

OpenStudy (n00bstyle):

You will see that when you divide e ( or whatever number) by infinity, you will end up having zero

OpenStudy (n00bstyle):

so: \[e^{-infinity}=\frac{ e }{ infinity }\rightarrow0\]

OpenStudy (n00bstyle):

And since the other part of your equation (in your question) is a constant (=1), you'll end up with 1 as an answer, because this doesn't include a variable...so this number just stays 1

OpenStudy (anonymous):

so the limit is 1?

OpenStudy (n00bstyle):

yes

OpenStudy (n00bstyle):

But the most important thing is the reformulation of e^(-infinity) = 1/(infinity) = 0 !!

OpenStudy (n00bstyle):

always try to fill in numbers like -infinity, 0, 1, and +infinity. It really helps you to 'see' or understand what happens to a graph

OpenStudy (n00bstyle):

For example: put in -infinity, do you understand what happens?

OpenStudy (n00bstyle):

\[e^{-(-infinity)}=e^{+infinity}=infinity\] So your graph will rise skyhigh towards infinity (for -infinity, so you go to the left of your graph)

OpenStudy (n00bstyle):

in your question I mean in this example

OpenStudy (anonymous):

okay i see what you're saying... and because it's negative that means that the entire function is decreasing right?

OpenStudy (n00bstyle):

yes! (ofcourse: THAT PART of the equation)

OpenStudy (n00bstyle):

The 1 still remains 1, because that part doesn't include a variable t

OpenStudy (n00bstyle):

Good luck

OpenStudy (anonymous):

thank you!!